OpenStudy (anonymous):

a bag of lollipops contain 20 cherry, 15 lemon, and 5 grape. If you close your eyes and take one lollipop from the bag, what is the probability of it being lime? (Express your answer as a fraction in simplest terms.)

OpenStudy (anonymous):

There is no lime in your bag...

OpenStudy (anonymous):

so ur chances are nun...now i want a limelollipop...nd u have nun!

OpenStudy (anonymous):

tease :(

OpenStudy (anonymous):

total possibilities for a lollipop=20+15+5=40 possibility of lime =15 p(lime)=total possibilities/possibility of lime =15/40 =3/8

OpenStudy (anonymous):

hoddo u get tht if there isnt any lime lollipops in his bag....

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

so am i shanna care to explain?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

total possibilities=total possible outcomes=total lollipops=20+15+5=40 we want probability of limes so total possibilities of lime =total possible outcomes of lime =total limes = 15 probability(lime)=total possibilities/possibility of lime =15/40 =3/8

OpenStudy (anonymous):

hope it helps!

OpenStudy (anonymous):

but that is lemon (15) not lime